OpenStudy (anonymous):

Evaluate the definite integral from x=0 to x=1 the cubed root of (1 + 7x) dx

OpenStudy (anonymous):

A substitution step would simplify the problem: \[\int\limits_{0}^{1}(1+7x)^{\frac{1}{3}}dx\] \[u = 1+7x\] \[\frac{du}{dx} = 7 \] \[dx = \frac{du}{7}\] \[\frac{1}{7}\int\limits_{}^{}u^{\frac{1}{3}}du\]

OpenStudy (anonymous):

you'll have to redo the bounds of integration (ie it was 0-1 when the integral was in terms of 'x', but now the integral is in terms of 'u')

OpenStudy (anonymous):

ohhh that was why i messed up. thank you!

OpenStudy (anonymous):

you got the new bounds?

OpenStudy (anonymous):

yea i thought so, but my answer was wrong. it's okay though. it's only one point off my homework. thanks anyway!